A ball of mass52. 4g is thrown from the ground into the air with an initial velocity of 16.3m/s at an angle of 27.4 with the horizontal. What are the values of its kinetic energy initially and just before it strikes the ground 

Grade:12th pass

1 Answers

Umar Farooq
5 years ago
Its initial velocity u=16.3 (cos 27.4) i +16.3 (sin 27.4)
Initial kinetic energy KEinitial=0.5 mu2 = 0.5 x 0.0524 x 16.3= 69.6 J
As there’s no force along horizontal direction vx=16.3 (cos 27.4)=14.47
vy= u (sin 27.4) + gT
T= 2u sin (27.4) /g
vy3u (sin 27.4)
v= 14.47 i + 22.5 j
v= 26.75
KEfinal =0.5 mv2 =0.5 x 0.0524 x 26.752=187.48 J